2. Using Table 3-3, please provide the calculations below by assuming: (a) cholera was a chronic illness; (b) the reported numbers for each year are newly diagnosed cases; and (c) that no death occurred during that time in the population of 1000. • Calculate incidence for cholera in 2013 • Calculate prevalence for cholera in 2015

TABLE 3-3 Cases of Selected Infrequently Reported Notifiable Diseases (< 1,000 Cases Reported)—United States, 2011–2015

Total Cases Reported by Year
Disease 2015* 2014 2013 2012 2011

Not reportable in all states.

*Case counts for 2015 are provisional.

Modified data from Centers for Disease Control and Prevention. Notifiable diseases and mortality tables. MMWR. 2016;65(24):ND-417.

Botulism, foodborne 37 15 4 27 24
Cholera 2 5 14 17 40
Hansen’s disease (leprosy)† 89 88 81 82 82
Rabies, human 1 1 2 1 6
Answer #1

Incidence = Total number of new cases of a particular disease during a given period   x 1000 or 100000

Total population at risk during the same time period

usually for infrequent disease larger value multiplied. so 100000 has taken

Incidence for cholera in 2013

= (14/ 1000)* 100000

= 0.014 cases per 100000

Prevalence = All new and old cases of a specific disease during a given time period x 100

Estimated population at risk

Prevalence for cholera in 2015

= (2+5+14+17+40/ 1000) x 100

=(78/1000) x 100

= 0.078 cases per 100